OpenStudy (anonymous):

Solve 11.3^(2x – 1) = 15.7 x ≈ 1.12 x ≈ 0.57 x ≈ 0.94 x ≈ 1.07

OpenStudy (callisto):

Similar to your previous question. Take log on both sides 11.3^(2x – 1) = 15.7 log [11.3^(2x – 1)] = log15.7 Use log property: \[loga^b = b\log a\] (2x-1) log 11.3 = log15.7 Divide both sides by log 11.3. What do you get so far?

OpenStudy (anonymous):

taking log of both sides. \[\Large \log(3)^{2x-1}=\log(15.7)\] or using the log property \[\Large \log(a)^n=nlog(a)\] \[\Large \implies (2x-1)\log(11.3)=\log(15.7)\] can you solve this now ?

OpenStudy (anonymous):

(2x-1)1.05=1.19 idk what to do with 2x-1

OpenStudy (anonymous):

2x-1=0.882

OpenStudy (anonymous):

0.94?

OpenStudy (callisto):

Not really.. 1.19/1.05 ≠ 0.882